A gear drive consists of two gears, A and B, and has a velocity ratio of 1.50. Gear A, the


smaller of the two gears, revolves at 126 rpm in the clockwise direction, and has 28 teeth.


If the gears have a module of 2 mm, determine:


i)


The number of teeth on Gear B.


ii)


The pitch (reference) diameters for the two gears.


iii) The addendum.


iv) The dedendum.


v)


The circular pitch.


vi) The tooth thickness.


vii) The speed of Gear B.


viii) The theoretical centre distance of the two gears.

1
Expert's answer
2022-02-06T11:34:09-0500

i)

VR=TB/TA

TB=VR*TA=1.5*28

TB=42

ii)

dA=mTA=2*28=56mm

dB=mTB=2*42=84mm

iii)

a=m=2mm

iv)

d=1,157*2=2.314mm

v)

p="\pi" m="\pi*2=6.28mm"

vi)

"t=\frac{\pi m}{2}=\frac{\pi*2}{2}=3.14mm"

vii)

VR="\frac{N_A}{N_B}"

"N_B=\frac{N_A}{VR}=\frac{126}{1.5}=84rpm"

viii)

"C=\frac{m(T_A+T_B)}{2}=\frac{2*(28+42)}{2}=70mm"